Establish a relationship among Young’s modulus (Y), bulk modulus (K), rigidity modulus (η), and Poisson’s ratio (σ).


1
Expert's answer
2021-03-10T17:20:23-0500
"\frac{3}{Y}=\frac{1}{\eta}+\frac{1}{3K}\\2\sigma=1-\frac{Y}{3K}"